OpenStudy (anonymous):

How would you transform the polar equation "r csc theta = 2" to an equation in rectangular coordinates? Answer needs to have an x and a y on one side. Equations to use: x = r cos theta; y = r sin theta; x^2 + y^2 = r^2

OpenStudy (dumbcow):

csc = 1/sin rewrite equation \[r = 2 \sin \theta\] square both sides to get an r^2 term \[r^2 = 4 \sin^2 \theta\] substitute ---> r^2 = x^2+y^2 and sin = y/r \[x^2 + y^2 = 4(\frac{y^2}{r^2})\] substitute for r^2 again \[x^2 +y^2 = \frac{4y^2}{x^2 + y^2}\]

OpenStudy (dumbcow):

wait sorry we can simply it further \[(x^2 +y^2)^2 = 4y^2\] sqrt both sides \[x^2 +y^2 = 2y\] now you have a circle

OpenStudy (anonymous):

Thank you so much for the response, unfortunately that is not the answer my teacher was looking for. The correct answer was \[x^2+(y-1)^2=1\] just in case somebody out there gets a similar question.

OpenStudy (dumbcow):

same thing just in different forms \[x^2 +y^2 = 2y\] \[x^2 + (y^2 -2y) = 0\] \[x^2 + (y^2 -2y +1) = 1\] \[x^2 + (y-1)^2 = 1\]